OpenStudy (anonymous):

How many solutions are there to the following system of equations? 3x + 7y = –10 15x + 35y = –45 0 1 2 infinitely many

OpenStudy (anonymous):

0. the equations are inconsistent.

OpenStudy (anonymous):

15x + 35y = -45 reduces to 3x + 7y = - 9 so we have : 3x + 7y = -9 --->(-1) 3x + 7y = -10 ------------- -3x - 7y = 9 (result of multiplying by -1) 3x + 7y = -10 -------------add 0 = -1 (incorrect) This system of equations has NO SOLUTIONS...or 0

OpenStudy (anonymous):

thanks that was conformation